Given 25 ml of a KCl solution with a concentration of 0.38 M and 18 ml of a solution of CaCl2 with a concentration of 0.19 M, what amount of each solution you should use to make 40 ml of solution with [Cl-]=0.30 M ?
Answer #1

KCl solution:

Using the releation: M * V = constant

Volume of KCl needed * Molarity of KCl = Volume of Cl- solution * molarity

V * 0.38 = 40 * 0.3

Solving, we get: V = 10.53 ml

CaCl2 Solution:

Since each CaCl2 has 2 chloride ions, the relation is:

2 * Volume of CaCl2 needed * Molarity of CaCl2 = Volume of Cl- solution * molarity

2 * V * 0.38 = 40 * 0.30

Solving, we get: V = 5.26 mL